OpenStudy (anonymous):

Help: y= (x^3 + 2x -4)^3 (4x-1)^5 Product rule?

OpenStudy (anonymous):

product rule + chain rule

OpenStudy (anonymous):

\[(fg)'=f'g+g'f\] \[f(x)=(x^3+2x-4)^3\] \[f'(x)=4(x^3+2x-4)^3(3x^2+2)\] \[g(x)=(4x-1)^5\] \[g'(x)=5(4x-1)^4\times 4\] i will let you put it together yourself

OpenStudy (anonymous):

sorry i made a mistake

OpenStudy (anonymous):

\[f(x)=(x^3+2x-4)^3\] \[f'(x)=3(x^2+2x-4)^2(3x^2+2)\]

OpenStudy (anonymous):

rest is correct though

OpenStudy (anonymous):

How do u know he wants to diff it?